Good day. I have 9 equations that I need help solving. Thank you in advance.
1. x/2=-16 x/2 is a fraction.
2. Fractions: 3/8x-1/3x+6=1/6x
3. fractions: -1/4(x-8)+1/2(x-2)=x+9
4. Solve for k: F=kn
5. Solve for a: P=a+b+c
6. Solve the inequality, write in interval notation. 3x+5+x<2+2x+3
7. same as 6: -x+1+6x _> -2+2x+24
8. 11x+2>2(4x+1)-x+7
9. The length of a rectangular boat storage shed must be 4 meters more than the width, and the perimeter must be at least 108 meters. What is the range of values for the width?

Answer by Alan3354(69443)

1. x/2=-16 x/2 is a fraction.
No kidding.
Multiply both sides by 2.
--------------------------------
What sort of help do you need? What have you tried?
Repost, showing your efforts.
------------------------------
2. Fractions: 3/8x-1/3x+6=1/6x
3. fractions: -1/4(x-8)+1/2(x-2)=x+9
4. Solve for k: F=kn
5. Solve for a: P=a+b+c
6. Solve the inequality, write in interval notation. 3x+5+x<2+2x+3
7. same as 6: -x+1+6x _> -2+2x+24
8. 11x+2>2(4x+1)-x+7
9. The length of a rectangular boat storage shed must be 4 meters more than the width, and the perimeter must be at least 108 meters. What is the range of values for the width?